# Fractal Patterns May Unravel the Intelligence in Next-Token Prediction

### v1

@@ 2.2 Self-similarity exponent
An integral process is said to be self-similar if it exhibits statistical self-similarity. More precisely, $(X_{t})_{t\in\mathbb{N}}$ is self-similar if $(X_{\tau t})_{t\in\mathbb{N}}$ is distributionally equivalent to $(\tau^{S}X_{t})_{t\in\mathbb{N}}$ for some exponent $\mathrm{S}$. Thus, scaling of time is equivalent to an appropriate scaling of space. We will refer to $\tau$ as the granularity level and to the exponent $\mathrm{S}$ as the self-similarity exponent. It is worth noting that $\mathrm{S}$ is also called the Hölder exponent (Watkins, 2019). Many time series in nature exhibit self-similar structures, such as human blood pressure and heart rate (Goldberger et al., 2002).
One approach for calculating the self-similarity exponent $\mathrm{S}$ is as follows. First, fix $\epsilon\ll 1$ and denote the $\tau$-increments by $(X_{t+\tau}-X_{t})_{t\in\mathbb{N}}$. These would correspond, for instance, to the number of bits used for clauses, sentences, paragraphs and longer texts as $\tau$ increases. In terms of the increment process $(x_{t})_{t\in\mathbb{N}}$, this corresponds to aggregating increments into “bursts”. Let $p_{\epsilon}(\tau)$ be the probability mass of the event $\{|X_{t+\tau}-X_{t}|\leq\epsilon\}_{t\in\mathbb{N}}$. Then, $\mathrm{S}$ can be estimated by fitting a power law relation $p_{\epsilon}(\tau)\sim\tau^{-S}$ (Watkins, 2019).
Figure 2 (top) plots the probability $p_{\epsilon}(\tau)$ against $\tau$ when $\epsilon=5\times 10^{-3}$ using PaLM2-L. We indeed observe a power law relation; i.e. linear in a log-log scale, with a median self-similarity exponent of $\mathrm{S}=0.59\pm 0.08$. Section 3 shows that the median $\mathrm{S}$ is robust to the choice of the LLM.
[…]
@@ 2.3 Hurst parameter
The Hurst parameter $\mathrm{H}\in[0,1]$ quantifies the degree of predictability or dependence over time (Hurst, 1951). It is calculated using the so-called rescaled-range (R/S) analysis. Let $(x_{t})_{t\in\mathbb{N}}$ be an increment process. For each $n\in\mathbb{N}$, write $y_{t}=x_{t}-\frac{1}{t}\sum_{k=0}^{t}x_{k}$ and $Y_{t}=\sum_{k=0}^{t}y_{t}$. The range and scale are defined, respectively, as $R(n)=\max_{t\leq n}Y_{t}-\min_{t\leq n}Y_{t}$ and $S(n)=\sigma\left(\{x_{k}\}_{k\leq n}\right)$, where $\sigma$ is the standard deviation. Then, the Hurst parameter $\mathrm{H}$ is estimated by fitting a power law relation $R(n)/S(n)\sim n^{\mathrm{H}}$. As stated earlier, for completely random processes, such as a simple Brownian motion, it can be shown that $\mathrm{H}=1/2$. In addition, $H>1/2$ implies dependence over time (Crovella & Bestavros, 1995; Willinger et al., 1995; Aref, 1998).
Writing $\rho_{n}=\mathbb{E}[(x_{t+n}x_{t}]$ for the autocovariance function of the increment process $(x_{t})_{t\in\mathbb{N}}$, the Hurst parameter satisfies $\mathrm{H}=1-\beta/2$ when $\rho_{n}\sim n^{-\beta}$ as $n\to\infty$ (Gneiting & Schlather, 2004; Crovella & Bestavros, 1995). Since in self-similar processes, $\mathrm{H}>1/2$ implies long-range dependence (LRD), LRD is equivalent to the condition that the autocovariances are not summable. In terms of the integral process, it can be shown that (Samorodnitsky, 2006):
$\lim_{n\to\infty}\frac{\mathrm{Var}(X_{n})}{n}=1+2\sum_{i=1}^{\infty}\rho_{i}.$ (1)
Hence, if $\mathrm{H}<1/2$, the auto-covariances are summable and $\mathrm{Var}(X_{n})$ grows, at most, linearly fast on $n$. On the other hand, if the process has LRD, $\mathrm{Var}(X_{n})$ grows superlinearly on $n$. In particular, using the Euler-Maclaurin summation formula (Apostol, 1999; Alabdulmohsin, 2018), one obtains $\mathrm{Var}(X_{n})\sim n^{2H}$ if $H>1/2$. Figure 4 plots the rescaled range $R(n)/S(n)$ against $n$. We observe a power law relation with a median Hurst parameter of $\mathrm{H}=0.70\pm 0.09$.
@@
$\mathrm{S}$
$0.53\pm.05$
$0.60\pm.05$
$0.61\pm.05$
$0.56\pm.03$
$0.62\pm.02$
$0.60\pm.07$
$0.42\pm.03$
$0.70\pm.03$
$\mathrm{H}$
$0.68\pm.01$
$0.79\pm.01$
$0.68\pm.00$
$0.70\pm.00$
$0.74\pm.01$
$0.65\pm.00$
$0.50\pm.01$
$0.72\pm.01$
$\mathrm{J}$
$0.46\pm.01$
$0.49\pm.00$
$0.49\pm.00$
$0.50\pm.00$
$0.52\pm.00$
0.44±.00plus-or-minus0.44.000.$4$00
$0.28\pm.00$
$0.49\pm.00$
T5-Decoder PaLM PaLM2
110M
340M
1B
5B
8B
62B
540B
XXS
XS
S
M
L
$\mathrm{S}$
.$58\pm.06$
.$60\pm.06$
.$60\pm.05$
.$58\pm.08$
.$60\pm.07$
.$62\pm.08$
.$64\pm.08$
.$59\pm.06$
.$57\pm.08$
.$56\pm.05$
.$59\pm.07$
.$60\pm.08$
$\mathrm{H}$
.$64\pm.08$
.$64\pm.08$
.$64\pm.09$
.$64\pm.08$
.66±.07plus-or-minus.66.07.$6$07
.$68\pm.07$
.$68\pm.07$
.66±.07plus-or-minus.66.07.$6$07
.66±.07plus-or-minus.66.07.$6$07
.$67\pm.08$
.$68\pm.09$
.$69\pm.09$
$\mathrm{J}$
.44±.06plus-or-minus.44.06.$4$06
.44±.06plus-or-minus.44.06.$4$06
.44±.06plus-or-minus.44.06.$4$06
.44±.06plus-or-minus.44.06.$4$06
.$47\pm.06$
.$47\pm.06$
.$48\pm.06$
.$47\pm.06$
.$47\pm.06$
.$48\pm.07$
.$48\pm.07$
.$49\pm.08$
@@ 2.4 Fractal dimension
Broadly speaking, the fractal dimension of an object describes its local complexity. For a geometric object $Z$, such as the Koch curve, let $\tau$ be a chosen scale (e.g. a short ruler for measuring lengths or a small square for areas). Let $N(\tau)$ be the minimum number of objects of scale $\tau$ that cover $Z$. Then, the fractal dimension of $Z$, also called its Hausdorff dimension, is: $\mathrm{D}=-\lim_{\tau\to 0}\left\{\frac{\log N(\tau)}{\log\tau}\right\}$ (Pilgrim & Taylor, 2018). For example, a line has a fractal dimension $1$, in agreement with its topological dimension, because $N(\tau)=C/\tau$ for some constant $C>0$.
By convention, an object is referred to as “fractal” if $\mathrm{D}$ is different from its topological dimension. For example, the fractal dimension of the Koch curve is about 1.26 when its topological dimension is 1. Fractals explain some puzzling observations, such as why estimates of the length of the coast of Britain varied significantly from one study to another, because lengths in fractals are scale-sensitive. Mandelbrot estimated the fractal dimension of the coast of Britain to be 1.25 (Mandelbrot, 1967).
The definition above for the fractal dimension $\mathrm{D}$ applies to geometric shapes, but an analogous definition has been introduced for stochastic processes. Let $(x_{t})_{t\in\mathbb{R}}$ be a stationary process with autocovariance $\rho_{n}$. Then, its fractal dimension $\mathrm{D}$ is determined according to the local behavior of $\rho_{n}$ at the vicinity of $n=0$, by first normalizing $(x_{t})_{t\in\mathbb{R}}$ to have a zero-mean and a unit variance, and modeling $\rho_{n}$ using a power law $\rho_{n}\sim 1-n^{\alpha}$ as $n\to 0^{+}$, for $\alpha\in(0,2]$. Then, the fractal dimension D[1, $\mathrm{D}\in[1,\,2]$ of $(x_{t})_{t\in\mathbb{R}}$ is defined by $\mathrm{D}=2-\alpha/2$ (Gneiting & Schlather, 2004). A value $\mathrm{D}\gg 1$ indicates a significant fractal structure.
It can be shown that $\mathrm{D}=2-\mathrm{S}$, where $\mathrm{S}$ is the self-similarity exponent (Gneiting & Schlather, 2004). For language, this gives a median fractal dimension of $\mathrm{D}=1.41\pm 0.08$.
@@ 2.5 Joseph effect
Next, we examine another related parameter that is commonly studied in self-similar processes. The motivation behind it comes from the fact that in processes with LRD, one often observes burstiness as shown in Figure 1; i.e. clusters over time in which the process fully resides on one side of the mean, before switching to the other. This is quite unlike random noise, for instance, where measurements are evenly distributed on both sides of the mean. The effect is often referred to as the Joseph effect, named after the biblical story of the seven fat years and seven lean years (Willinger et al., 1995; Mandelbrot & Wallis, 1968; Watkins, 2019).
A common way to quantify the Joseph effect for integral processes $(X_{t})_{t\in\mathbb{N}}$ is as follows (Watkins, 2019). First, let $\sigma_{\tau}$ be the standard deviation of the $\tau$-increments $X_{t+\tau}-X_{t}$. Then, fit a power law relation $\sigma_{\tau}\sim\tau^{\mathrm{J}}$. The exponent $\mathrm{J}$ here is called the Joseph exponent. In an idealized fractional Brownian motion, both $\mathrm{J}$ and the self-similarity exponent $\mathrm{S}$ coincide. Figure 5 provides the detailed empirical results. Overall, we obtain an estimate of $\mathrm{J}=0.49\pm 0.08$, which is intriguing because $J=0.5$ corresponds to self-similar processes with independent increments.
@@ 3 Analysis
Table 1 compares the estimated fractal parameters across different domains, such as ArXiv, Github and Wikipedia. In general, most domains share similar self-similarity and Hurst exponents with a few exceptions. The first notable exception is DM-Mathematics, which has a Hurst parameter of about 0.5. To recall, a value of $\mathrm{H}=0.5$ indicates that the data does not exhibit long-range dependence (LRD). Upon closer inspection, however, a value of $\mathrm{H}=0.5$ is not surprising for DM-Mathematics because its documents consist of independent mathematical questions as shown in Figure 7. The second notable observation is the relatively larger value of $\mathrm{H}=0.79$ in GitHub, indicating more structure in code. This is in agreement with earlier findings by Kokol & Podgorelec (2000) who estimated LRD in computer languages to be greater than in nature language. In Table 2, we compare the three fractal parameters $\mathrm{S}$, $\mathrm{H}$ and $\mathrm{J}$ using different families of LLM and different model sizes. Overall, we observe that the estimated parameters are generally robust to the choice of the architecture.
Downstream Performance.
By definition, fractal parameters are calculated on the sequence of log-perplexity scores after normalizing them to zero-mean and unit variance. Hence, they may offer an assessment of downstream performance that improves upon using a perplexity-based metric like bits-per-byte (BPB) alone.
To test this hypothesis, we evaluate the 12 models in Table 2 on challenging downstream zero- and few-shot benchmarks focusing on language understanding and reasoning. We include results for 0-shot (0S) and 3-shot (3S) evaluation for BIG-Bench Hard tasks (Srivastava et al., 2022; Suzgun et al., 2022) reporting both direct and chain-of-thought (CoT) prompting results following Chung et al. (2022). In addition we report 0-shot and 5-shot (5S) MMLU (Hendrycks et al., 2020), and 8-shot (8S) GSM8K (Cobbe et al., 2021) with CoT. Raw accuracy is reported for all tasks. BBH and MMLU scores are averaged across all 21 tasks and 57 subjects, respectively. All prompt templates for our evaluation are taken from Chung et al. (2022); Longpre et al. (2023), which we refer the reader to for more details. We prompt all models using a 2048 context length. See Table 9 of Appendix C for the full results.
The first (surprising) observation is that the median Hurst parameter is itself strongly correlated with the BPB scores with an absolute Pearson correlation coefficient of 0.83, even though the Hurst exponent is calculated after normalizing all token losses to zero-mean and unit variance! Informally, this implies that second-order statistics on the sequence of token losses of a particular model can predict its mean! The self-similarity exponent, by contrast, has an absolute Pearson correlation of 0.23 with BPB.
Figure 6 displays downstream performance against both the median Hurst exponent and the median BPB score, where median values are calculated on the 8 domains in The Pile benchmark listed in Table 1. In general, both the BPB score and the median Hurst are good predictors of downstream performance. However, we observe that improvements in BPB alone without impacting the median Hurst exponent do not directly translate into improvements downstream. This is verified quantitatively in Table 3, which reports the adjusted $R^{2}$ values – the proportion of variance in each downstream metric that can be predicted using BPB, $\mathrm{H}$, or by combining them together into $\mathrm{H}_{B}=1/\mathrm{BPB}+\mathrm{H}$, with BPB replaced with its reciprocal so that higher values are better. We observe that $\mathrm{H}_{B}$ yields indeed a stronger predictor of downstream performance. For ranking, however, BPB alone is sufficient. See Appendix C for similar analysis using the exponents $\mathrm{S}$ and $\mathrm{J}$.
Table 3: Adjusted $R^{2}$, which measures the proportion of variation in downstream performance (row) predictable by a linear regressor with the given input (column). The combined metric $\mathrm{H}_{B}=1/\mathrm{BPB}+\mathrm{H}$ predicts downstream performance better in all downstream metrics, compared to BPB alone. $\mathrm{S}$ and $\mathrm{J}$ do not yield such improvements (see Appendix C). For ranking, we report Spearman correlations, which suggest that BPB is sufficient.
Magnitude Ranking
BPB
$\mathrm{H}$
$\mathrm{H}_{B}$
BPB
$\mathrm{H}_{B}$
0S BBH Direct
0.785
0.841
0.883
0.958
0.958
0S MMLU
0.653
0.831
0.825
0.769
0.769
0S BBH+MMLU
0.685
0.849
0.852
0.930
0.930
3S BBH Direct
0.767
0.895
0.926
1.000
1.000
3S BBH CoT
0.881
0.892
0.979
1.000
1.000
5S MMLU
0.660
0.853
0.832
0.783
0.783
8S GSM8K CoT
0.654
0.867
0.851
0.993
0.993
FS BBH+MMLU+GSM8K
0.717
0.890
0.891
1.000
1.000
Figure 6: Downstream metric, indicated by bubble size, is plotted vs. the median Hurst and the median BPB for all 12 language models.
Table 4: Downstream performance comparison for three decoder-only T5.1.1. models pretrained on 100B tokens with either 2K, 4K, or 8K context lengths.
2K
4K
8K
0S BBH Direct
01.81
01.68
01.76
0S MMLU
25.73
26.04
25.81
0S BBH+MMLU
13.39
13.49
13.42
3S BBH Direct
21.35
24.76
23.14
3S BBH CoT
16.87
12.21
07.14
5S MMLU
26.57
26.69
27.07
8S GSM8K CoT
01.06
01.21
01.74
FS BBH + MMLU+GSM8K
15.58
15.46
14.65
Context Length at Training Time.
Finally, self-similarity and long-range dependence point to an intriguing possibility: the importance of training the model with extensive contexts in order to capture the fractal-nature of language, which may elevate the model’s capabilities regardless of the context length needed during inference. To test this hypothesis, we pretrain three decoder-only T5.1.1 models with 1B parameters on SlimPajama-627B (Soboleva et al., 2023) for up to 100B tokens using three context lengths: 2K, 4K and 8K, all observing the same number of tokens per batch. We use SlimPajama-627B instead of C4 because most documents in C4 are short ($\approx 94\%$ of them are $<2K$ tokens in length). Refer to Appendix A for details. These models are, then, evaluated on the same downstream benchmarks listed in Figure 6 and Table 3. As shown in Table 4, however, we do not observe any improvements in performance with context length in this particular setup.
Figure 7: Two examples of documents from the DM-Mathematics subset of The Pile benchmark (Gao et al., 2020). Each document comprises of multiple independent questions. The lack of LRD in this data is reflected in its Hurst parameter of $\mathrm{H}=0.50\pm 0.01$
@@
4 Related Works

### v2

@@ 2.2 Self-similarity exponent
An integral process is said to be self-similar if it exhibits statistical self-similarity. More precisely, $(X_{t})_{t\in\mathbb{N}}$ is self-similar if $(X_{\tau t})_{t\in\mathbb{N}}$ is distributionally equivalent to $(\tau^{S}X_{t})_{t\in\mathbb{N}}$ for some exponent $\mathrm{S}$. Thus, scaling of time is equivalent to an appropriate scaling of space. We will refer to $\tau$ as the granularity level and to the exponent $\mathrm{S}$ as the self-similarity or Hölder exponent [64]. Many time series in nature exhibit self-similar structures, such as human blood pressure and heart rate [27].
One approach for calculating $\mathrm{S}$ is as follows. Fix $\epsilon\ll 1$ and denote the $\tau$-increments by $(X_{t+\tau}-X_{t})_{t\in\mathbb{N}}$. These would correspond, for instance, to the number of bits used for clauses, sentences, paragraphs and longer texts as $\tau$ increases. In terms of the increment process $(x_{t})_{t\in\mathbb{N}}$, this corresponds to aggregating increments into “bursts”. Let $p_{\epsilon}(\tau)$ be the probability mass of the event $\{|X_{t+\tau}-X_{t}|\leq\epsilon\}_{t\in\mathbb{N}}$. Then, $\mathrm{S}$ can be estimated by fitting a power law relation $p_{\epsilon}(\tau)\sim\tau^{-S}$ [64]. Generally, $\mathrm{S}$ is robust to the choice of $\epsilon\in[10^{-3},10^{-2}]$ as shown in Figure 4 (left) so we fix it to $\epsilon=5\times 10^{-3}$.
Figure 2 plots the probability $p_{\epsilon}(\tau)$ against $\tau$ using PaLM2-L. We indeed observe a power law relation; i.e. linear in a log-log scale, with a median self-similarity exponent of $\mathrm{S}=0.59\pm 0.08$. Section 3 shows that the median $\mathrm{S}$ is robust to the choice of the LLM.
@@ 2.3 Hurst parameter
The Hurst parameter $\mathrm{H}\in[0,1]$ quantifies the degree of predictability or dependence over time [31]. It is calculated using the so-called rescaled-range (R/S) analysis. Let $(x_{t})_{t\in\mathbb{N}}$ be an increment process. For each $n\in\mathbb{N}$, write $y_{t}=x_{t}-\frac{1}{t}\sum_{k=0}^{t}x_{k}$ and $Y_{t}=\sum_{k=0}^{t}y_{t}$. The range and scale are defined, respectively, as $R(n)=\max_{t\leq n}Y_{t}-\min_{t\leq n}Y_{t}$ and $S(n)=\sigma\left(\{x_{k}\}_{k\leq n}\right)$, where $\sigma$ is the standard deviation. Then, the Hurst parameter $\mathrm{H}$ is estimated by fitting a power law relation $R(n)/S(n)\sim n^{\mathrm{H}}$. As stated earlier, for completely random processes, such as a simple Brownian motion, it can be shown that $\mathrm{H}=1/2$. In addition, $H>1/2$ implies dependence over time [16, 65, 8].
Writing $\rho_{n}=\mathbb{E}[(x_{t+n}x_{t}]$ for the autocovariance function of the increment process $(x_{t})_{t\in\mathbb{N}}$, the Hurst parameter satisfies $\mathrm{H}=1-\beta/2$ when $\rho_{n}\sim n^{-\beta}$ as $n\to\infty$ [26, 16]. Since in self-similar processes, $\mathrm{H}>1/2$ implies long-range dependence (LRD), LRD is equivalent to the condition that the autocovariances are not summable. In terms of the integral process, it can be shown that [57]: $\lim_{n\to\infty}\frac{\mathrm{Var}(X_{n})}{n}=1+2\sum_{i=1}^{\infty}\rho_{i}$. Hence, if $\mathrm{H}<1/2$, the auto-covariances are summable and $\mathrm{Var}(X_{n})$ grows, at most, linearly fast on $n$. On the other hand, if the process has LRD, $\mathrm{Var}(X_{n})$ grows superlinearly on $n$. In particular, using the Euler-Maclaurin summation formula [7, 2], one obtains $\mathrm{Var}(X_{n})\sim n^{2H}$ if $H>1/2$. Figure 3 plots the rescaled range $R(n)/S(n)$ against $n$. We observe a power law relation with a median Hurst parameter of $\mathrm{H}=0.70\pm 0.09$.
@@
$\mathrm{S}$ $0.53\pm.05$ $0.60\pm.05$ $0.61\pm.05$ $0.56\pm.03$ $0.62\pm.02$ $0.60\pm.07$ $0.42\pm.03$ $0.70\pm.03$
$\mathrm{H}$ $0.68\pm.01$ $0.79\pm.01$ $0.68\pm.00$ $0.70\pm.00$ $0.74\pm.01$ $0.65\pm.00$ $0.50\pm.01$ $0.72\pm.01$
$\mathrm{J}$ $0.46\pm.01$ $0.49\pm.00$ $0.49\pm.00$ $0.50\pm.00$ $0.52\pm.00$ 0.44±.00plus-or-minus0.44.000.$4$00 $0.28\pm.00$ $0.49\pm.00$
@@ 2.4 Fractal dimension
Broadly speaking, the fractal dimension of an object describes its local complexity. For a geometric object $Z$, such as the Koch curve, let $\tau$ be a chosen scale (e.g. a short ruler for measuring lengths or a small square for areas). Let $N(\tau)$ be the minimum number of objects of scale $\tau$ that cover $Z$; i.e. contain it entirely. Then, the fractal dimension of $Z$, also called its Hausdorff dimension, is: $\mathrm{D}=-\lim_{\tau\to 0}\left\{\frac{\log N(\tau)}{\log\tau}\right\}$ [53]. For example, a line has a fractal dimension $1$, in agreement with its topological dimension, because $N(\tau)=C/\tau$ for some constant $C>0$.
By convention, an object is referred to as “fractal” if $\mathrm{D}$ is different from its topological dimension. For example, the fractal dimension of the Koch curve is about 1.26 when its topological dimension is 1. Fractals explain some puzzling observations, such as why estimates of the length of the coast of Britain varied significantly from one study to another, because lengths in fractals are scale-sensitive. Mandelbrot estimated the fractal dimension of the coast of Britain to be 1.25 [42].
The definition above for the fractal dimension $\mathrm{D}$ applies to geometric shapes, but an analogous definition has been introduced for stochastic processes. Let $(x_{t})_{t\in\mathbb{R}}$ be a stationary process with autocovariance $\rho_{n}$. Then, its fractal dimension $\mathrm{D}$ is determined according to the local behavior of $\rho_{n}$ at the vicinity of $n=0$, by first normalizing $(x_{t})_{t\in\mathbb{R}}$ to have a zero-mean and a unit variance, and modeling $\rho_{n}$ using a power law $\rho_{n}\sim 1-n^{\alpha}$ as $n\to 0^{+}$, for $\alpha\in(0,2]$. Then, the fractal dimension D[1, $\mathrm{D}\in[1,\,2]$ of $(x_{t})_{t\in\mathbb{R}}$ is defined by $\mathrm{D}=2-\alpha/2$ [26]. It can be shown that $\mathrm{D}=2-\mathrm{S}$ [26]. For language, this gives a median fractal dimension of $\mathrm{D}=1.41\pm 0.08$.
@@ 2.5 Joseph effect
Finally, we examine another related parameter that is commonly studied in self-similar processes. The motivation behind it comes from the fact that in processes with LRD, one often observes burstiness as shown in Figure 1; i.e. clusters over time in which the process fully resides on one side of the mean, before switching to the other. This is quite unlike random noise, for instance, where measurements are evenly distributed on both sides of the mean. The effect is often referred to as the Joseph effect, named after the biblical story of the seven fat years and seven lean years [65, 45, 64].
A common way to quantify the Joseph effect for integral processes $(X_{t})_{t\in\mathbb{N}}$ is as follows [64]. First, let $\sigma_{\tau}$ be the standard deviation of the $\tau$-increments $X_{t+\tau}-X_{t}$. Then, fit a power law relation $\sigma_{\tau}\sim\tau^{\mathrm{J}}$. The exponent $\mathrm{J}$ here is called the Joseph exponent. In an idealized fractional Brownian motion, both $\mathrm{J}$ and the self-similarity exponent $\mathrm{S}$ coincide. Figure 5 provides the detailed empirical results. Overall, we find that $\mathrm{J}=0.49\pm 0.08$.
@@ 3 Analysis
Table 1 compares fractal parameters across different domains, such as ArXiv, Github and Wikipedia. In general, most domains share similar self-similarity and Hurst exponents with a few exceptions. The first notable exception is DM-Mathematics, which has a Hurst parameter of about 0.5, indicating a lack of LRD. Upon closer inspection, however, a value of $\mathrm{H}=0.5$ is not surprising for DM-Mathematics because its documents consist of independent mathematical questions as shown in Figure 6. In Figure 4 (right), we plot the partial autocorrelation function for each of the 8 domains against time lag (context length). Indeed, we see that DM-Mathematics shows markedly less dependence compared to the other domains. The second notable observation is the relatively larger value of $\mathrm{H}=0.79$ in GitHub, indicating more structure in code. This is in agreement with earlier findings by Kokol and Podgorelec, [35] who estimated LRD in computer languages to be greater than in natural language. In Table 2, we compare the three fractal parameters $\mathrm{S}$, $\mathrm{H}$ and $\mathrm{J}$ using different families of LLM and different model sizes. Overall, we observe that the parameters are generally robust to the choice of the architecture.
T5-Decoder PaLM PaLM2
110M 340M 1B 5B 8B 62B 540B XXS XS S M L
Self-similarity exponent $\mathrm{S}$
.$58^{\pm.06}.$$60^{\pm.06}.$$60^{\pm.05}.$$58^{\pm.08}.$$60^{\pm.07}.$$62^{\pm.08}.$$64^{\pm.08}.$$59^{\pm.06}.$$57^{\pm.08}.$$56^{\pm.05}.$$59^{\pm.07}.$$60^{\pm.08}.$
Hurst exponent $\mathrm{H}$
.$64^{\pm.08}.$$64^{\pm.08}.$$64^{\pm.09}.$$64^{\pm.08}.$$66^{\pm.07}.$$68^{\pm.07}.$$68^{\pm.07}.$$66^{\pm.07}.$$66^{\pm.07}.$$67^{\pm.08}.$$68^{\pm.09}.$$69^{\pm.09}.$
Joseph exponent $\mathrm{J}$
.$44^{\pm.06}.$$44^{\pm.06}.$$44^{\pm.06}.$$44^{\pm.06}.$$47^{\pm.06}.$$47^{\pm.06}.$$48^{\pm.06}.$$47^{\pm.06}.$$47^{\pm.06}.$$48^{\pm.07}.$$48^{\pm.07}.$$49^{\pm.08}.$
Figure 6: Two examples of documents from the DM-Mathematics subset of The Pile benchmark [23]. Each document comprises of multiple independent questions. The lack of LRD in this data is reflected in its Hurst parameter of $\mathrm{H}=0.50\pm 0.01$
@@
Downstream Performance.
By definition, fractal parameters are calculated on the sequence of negative log-probability scores after normalizing them to zero-mean and unit variance. Hence, they may offer an assessment of downstream performance that improves upon using a perplexity-based metric like bits-per-byte (BPB) alone. To test this hypothesis, we evaluate the 12 models in Table 2 on challenging downstream zero- and few-shot benchmarks focusing on language understanding and reasoning. We include results for 0-shot (0S) and 3-shot (3S) evaluation for BIG-Bench Hard tasks [62, 63] reporting both direct and chain-of-thought (CoT) prompting results following Chung et al., [13]. In addition we report 0-shot and 5-shot (5S) MMLU [30], and 8-shot (8S) GSM8K [14] with CoT. Raw accuracy is reported for all tasks. BBH and MMLU scores are averaged across all 21 tasks and 57 subjects, respectively. All prompt templates for our evaluation are taken from Chung et al., [13], Longpre et al., [41], which we refer the reader to for more details. We prompt all models using a 2048 context length. See Table 8 of Appendix C for the full results.
The first (surprising) observation is that the median Hurst parameter is itself strongly correlated with the BPB scores with an absolute Pearson correlation coefficient of 0.83, even though the Hurst exponent is calculated after normalizing all token losses to zero-mean and unit variance! Informally, this implies that second-order statistics on the sequence of token losses of a particular model can predict its mean! Self-similarity exponent, by contrast, has an absolute correlation of 0.23 with BPB.
Figure 7 displays downstream performance against both the median Hurst exponent and the median BPB score, where median values are calculated on the 8 domains in The Pile benchmark listed in Table 1. In general, both the BPB score and the median Hurst are good predictors of downstream performance. However, we observe that improvements in BPB alone without impacting the median Hurst exponent do not directly translate into improvements downstream. This is verified quantitatively in Table 3 (middle), which reports the adjusted $R^{2}$ values – the proportion of variance in each downstream metric that can be predicted using BPB, $\mathrm{H}$, or by combining them together into $\mathrm{H}_{B}=1/\mathrm{BPB}+\mathrm{H}$, with BPB replaced with its reciprocal so that higher values are better. We observe that $\mathrm{H}_{B}$ yields indeed a stronger predictor of downstream performance. Hence, while $\mathrm{H}$ and BPB are correlated, combining them yields a better predictor, so each of $\mathrm{H}$ and BPB conveys useful information not captured by the other metric. See Appendix C for similar analysis using the exponents $\mathrm{S}$ and $\mathrm{J}$.
Table 3: middle three columns show the adjusted $R^{2}$: the proportion of variation in downstream performance (row) predictable by a linear function of the input (column). Median Hurst ($\mathrm{H}$) and (especially) the combined metric $\mathrm{H}_{B}$ predict downstream performance better than BPB alone. $\mathrm{S}$ and $\mathrm{J}$ do not give any improvement (see Appendix C). right: the downstream performance for three decoder-only T5.1.1. models pretrained on 100B tokens with 2K, 4K, or 8K context lengths.
Benchmark BPB $\mathrm{H}$ $\mathrm{H}_{B}$ 2K 4K 8K
0S BBH Direct 0.785 0.841 0.883 01.81 01.68 01.76
0S MMLU 0.653 0.831 0.825 25.73 26.04 25.81
0S BBH+MMLU 0.685 0.849 0.852 13.39 13.49 13.42
3S BBH Direct 0.767 0.895 0.926 21.35 24.76 23.14
3S BBH CoT 0.881 0.892 0.979 16.87 12.21 07.14
5S MMLU 0.660 0.853 0.832 26.57 26.69 27.07
8S GSM8K CoT 0.654 0.867 0.851 01.06 01.21 01.74
FS BBH+MMLU+GSM8K 0.717 0.890 0.891 15.58 15.46 14.65
Figure 7: Downstream metric, indicated by bubble size where larger is better, is plotted vs. the median Hurst and the median BPB for all 12 language models.
Context Length at Training Time (Negative Result).
Finally, we present a negative result. Self-similarity and LRD point to an intriguing possibility: the importance of training the model with extensive contexts in order to capture the fractal-nature of language, which may elevate the model’s capabilities regardless of the context length needed during inference. To test this hypothesis, we pretrain three decoder-only T5.1.1 models with 1B parameters on SlimPajama-627B [61] for up to 100B tokens using three context lengths: 2K, 4K and 8K, all observing the same number of tokens per batch. We use SlimPajama-627B instead of C4 because most documents in C4 are short ($\approx 94\%$ of them are $<2K$ tokens in length). Refer to Appendix A for details. These models are, then, evaluated on the same downstream benchmarks listed in Figure 7. As shown in Table 3 (right) however, we do not observe any improvements in performance with context length in this particular setup.
4 Related Works and Directions for Future Research